OpenStudy (anonymous):

how do you do these? I got 2,3, and 4. But I NEED #1 If you could give me all of them that'd be great

OpenStudy (anonymous):

http://www.mathbits.com/caching/PC358366.html

OpenStudy (anonymous):

so sin x=8/17. cos x=sqrt(1-xin^2 x)=sqrt(1-64/289)=sqrt(225/289)=-15/17. as it is given cos x is negative. so, sec x=1/cos x=-17/15

OpenStudy (anonymous):

have u figured it out yet mister?

OpenStudy (anonymous):

so the overall answer for #1 is ?

OpenStudy (anonymous):

Did you understand the method or not??

OpenStudy (anonymous):

not really

OpenStudy (anonymous):

Look all u have to do is go back to where that person explained it to you and try harder to under stand it!

OpenStudy (anonymous):

bye:D

OpenStudy (anonymous):

Do you know sin^2 x + cos ^2 x=1?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

I don't understand what the final answer is for #1. Is it (-17/15) ?

OpenStudy (anonymous):

Yeah.

OpenStudy (anonymous):

For the rest I have: 2) sin(arctan(1)) 3) 5.196152 4) 8.9644 are those right?

OpenStudy (anonymous):

#1 \[\cos \Theta\] =-15/17 \[\sec \Theta\] =-17/15 #2 \[\sin(\tan^{-1} (1)=\] \[\sin 45 =\] \[\sqrt{2}/2\] #3 \[6(1/\sqrt{3})-2(-\sqrt{3}/2)\] \[6(\sqrt{3}/3) + \sqrt{3}\] \[2\sqrt{3} + \sqrt{3}\] \[3\sqrt{3}\] #4 tan 35 = a/10 tan 62 = b/10 a = 0.7002*10 b=1.8807*10 a = 7.002 b = 18.807 Distance between towns is: 18.807-7.002 ans: 11.805

OpenStudy (anonymous):

So at the bottom to find the URL i keep doing it but it won't work